OpenStudy (anonymous):

precalc: Find the vertices of the hyperbola with equation: ((x+2)^2/4) - ((y+1)^2/16) = 1

OpenStudy (anonymous):

still dont understand what the vertices are....

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

center is at \[(-2,-1)\]

OpenStudy (anonymous):

since you have \[\frac{(x+2)^2}{4}\] and \[\sqrt{4}=2\] they are two units to the right and left of the center. hence they are at \[(-4,-1),(0,-1)\]